OpenStudy (anonymous):

REDUCE THE FRACTION 20/100

OpenStudy (anonymous):

1/5

OpenStudy (shadowlegendx):

shanerocks it is against he CoC to give out answers without any effort shown

OpenStudy (shadowlegendx):

As in no work being shown or any formulas

OpenStudy (anonymous):

o srry

OpenStudy (anonymous):

\[\frac{ 20 }{ 100 }=\frac{ 10 }{ 50 }=\frac{ 5 }{ 25 }=\frac{ 1 }{ 5 }\] you must divide by 2, then 2 again, and finally divide by 5

OpenStudy (♪chibiterasu):

They both have a common factor of 20, so divide both the numerator and denominator by 20.

OpenStudy (anonymous):

Translate the following written statement into an equation. Five plus twice a number is thirty-one. can u help me with this one

OpenStudy (♪chibiterasu):

Shanerocks, please ask that in your own separate question.

OpenStudy (anonymous):

ok................ 5 + 2n = 30 because twice a number is 2 * a number and +5 = 30 so thats makes the equation 5 + 2n = 30 @shanerocks